two years ago a father was five times as old as his son.two years later his age will be 8 more than 3 times the age of the son. find the present ages of father and son

Answer by Cromlix(4381)

Father = 5 * S
F = 5S
F = 5S
Two years later
F + 2 = 8 + 3(S + 2)
Replace F with 5S
5S + 2 = 8 + 3(S + 2)
5S + 2 = 8 + 3S + 6
5S - 3S = 8 + 6 - 2
2S = 12
S = 6
Present ages:
Son = 8 yrs old.
Father = 32 years old.
Hope this helps.
:-)